### v3

@@ 3.2 Projection of the image features into the pre-shape space
According to the theory of shape space, the vertices of a shape structure are typically required to be two-dimensional or three-dimensional. Therefore, the features obtained from the ViT network need to be up-dimensioned to twice their original number before being projected into a pre-shape space and a shape space. The up-dimensioned step can be achieved by adding additional dimensions or using techniques such as interpolation or transformation to match the required dimensional space of the shape structure.
In this paper, the up-dimensioned method is to copy a feature into two identical features. The image feature is assumed to be represented as $V_{0}=\{x_{1},x_{2},...,x_{n}\}$, where $n$ is the dimension of the feature. Thus, the up-dimensioned function $f(x)=x$, i.e., $V_{0}$ is transformed to $V_{1}=\{x_{1},y_{1},x_{2},y_{2},...,x_{n},y_{n}|y_{i}=x_{i},i\in[1,n]\}$. To mitigate the effect of the feature positions, the means of the features are removed, and the new feature $V_{2}=\{x_{1}^{{}^{\prime}},y_{1}^{{}^{\prime}},x_{2}^{{}^{\prime}},y_{2}^{{}^{\prime}},...,x_{n}^{{}^{\prime}},y_{n}^{{}^{\prime}}|x_{i}^{{}^{\prime}}=x_{i}-X_{\mu},y_{i}^{{}^{\prime}}=y_{i}-Y_{\mu},i\in[1,n]\}$, where $X_{\mu},Y_{\mu}$ are the means of $\{x_{i}\}$ and $\{y_{i}\}$, respectively. Then, the scaling of the feature $V_{2}$ is excluded with formula $z=\frac{V_{2}}{||V_{2}||}$, where $||V_{2}||$ means the Euclidean norm of $V_{2}$. Thus, the image feature $\frac{V_{2}}{||V_{2}||}$ is projected into a pre-shape space as $z$.
3.3 Construction of a Geodesic curve to fit the features
To construct a Geodesic curve in the pre-shape space, more than two image samples are required. Different Geodesic curves represent different categories or classes of images. As mentioned in the above subsections, the features are extracted from the multiple images with the ViT network model, and projected into a pre-shape space, denoted as $Z=\{z_{1}^{1},...,z_{1}^{M_{1}},...,z_{N}^{1},...,z_{N}^{M_{N}}\}$, where $N$ represents the total number of image categories, $M_{i}$ represents the number of images in the $i$-th category, and $z_{i}^{j}$ denotes $j$-th sample of the $i$-th category.
The set of the Geodesic curves representing multiple image categories in the pre-shape space is denoted as $\{\Gamma_{(v_{i}^{*},w_{i}^{*})}|i\in(1,N)\}$, where $v_{i}^{*}$ and $w_{i}^{*}$ represents the optimal parameters of the Geodesic curve, as $v_{1}$ and $v_{2}$ in Formula (5). To obtain the optimal $v_{i}^{*},w_{i}^{*}$ of the Geodesic curve $\Gamma_{(v_{i}^{*},w_{i}^{*})}$, the first step is to calculate the Geodesic distance among all features. Then, the feature with the maximum sum of distances from other features is selected as the initialization parameter $v_{i}^{*}$, defined as follows:
@@ 3.3 Construction of a Geodesic curve to fit the features
where $d(z_{i}^{k},z_{i}^{j})$ means the Geodesic distance between $z_{i}^{k}$ and $z_{i}^{j}$, calculated with Formula (2). Subsequently, $w_{i}^{*}$ is obtained through Geodesic curve $\Gamma_{(w_{i}^{0},w_{i}^{1})}(\cdot)$, which is determined by the two features $w_{i}^{0}$ and $w_{i}^{1}$ furthest from $v_{i}^{*}$. As shown in Figure 3 (c), the two features, $w_{i}^{0}$ and $w_{i}^{1}$, corresponding to the maximum two Geodesic distance from $v_{i}^{*}$, are calculated using the following formulas:
@@
where $Q_{i}=\{z_{i}^{1},z_{i}^{2},...,z_{i}^{M_{i}}\}$ to represent the set of features for the $i$-th category $\{z_{i}^{1},z_{i}^{2},...,z_{i}^{M_{i}}\}$. $Q_{i}-\{\#\}$ means if the features ##\## are in $Q_{i}$, they are removed from $Q_{i}$.
Then, a Geodesic curve $\Gamma_{(w_{i}^{0},w_{i}^{1})}(\cdot)$ between $w_{i}^{0}$ and $w_{i}^{1}$ is calculated with Formula (5). A set of features $\hat{w}_{i}^{\{1,...,S\}}$ is obtained by following the curve $\Gamma_{(w_{i}^{0},w_{i}^{1})}(\cdot)$, denoted as $\{\Gamma_{(w_{i}^{0},w_{i}^{1})}(s_{i}^{1}),\Gamma_{(w_{i}^{0},w_{i}^{1})}(s_{i}^{2}),...,\Gamma_{(w_{i}^{0},w_{i}^{1})}(s_{i}^{S})\}$, where $S$ represents the number of temporary features on $\Gamma_{(w_{i}^{0},w_{i}^{1})}(\cdot)$. Finally, among $v_{i}^{*}$ and $\hat{w}_{i}^{\{1,...,S\}}$, several Geodesic curves $\Gamma_{(v_{i}^{*},\hat{w}_{i}^{\{1,...,S\}})}(\cdot)$ are described. The sum of squared Geodesic distance is calculated from the $i$-th category of features $z_{i}^{1},z_{i}^{2},...,z_{i}^{M_{i}}$ to $\Gamma_{(v_{i}^{*},\hat{w}_{i}^{\{1,...,S\}})}(\cdot)$. The features $\hat{w}_{i}^{*}$ corresponding to the shortest sum is calculated as follows:
@@
and the step is illustrated in Figure 3 (f) and Figure 3 (g). Subsequently, assign the variable $w_{i}^{*}$ to $v_{i}^{*}$ and calculate the two furthest features from $v_{i}^{*}$, as illustrated in Figure 3 (h). Repeat the steps for obtaining $w_{i}^{*}$ as previously described in Figures 3 (i) to (j) until $w_{i}^{*}$ and $v_{i}^{*}$ cease to change. Consequently, $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$ denotes the optimal Geodesic curve fitting the distribution of feature $z_{i}^{*}$, as depicted in Figure 3 (k). The detailed procedure for FAGC of features for a specific category is outlined in Algorithm  1.
All Geodesic curves for different categories, denoted as $\{\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)|i\in\{1,N\}\}$, can be constructed with the above mentioned method. And, with utilizing Formula (2), the Geodesic distance $d(v_{i}^{*},w_{i}^{*})$ between all pairs of $v_{i}^{*}$ and $w_{i}^{*}$ is calculated. The distance indicates the range of the angle value $s$ within the Geodesic curve function $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$. A uniform distribution $s\sim U(0,d(v_{i}^{*},w_{i}^{*}))$ is employed to sample a set of angle values $\{s_{i}^{j}|j\in\{1,K\}\}$, where $K$ represents the number of sampled angle values for the Geodesic curves of the $i$-th category. The new features $Z^{\prime}=\{z_{1}^{{}^{\prime}1},...,z_{1}^{{}^{\prime}K},...,z_{N}^{{}^{\prime}1},...,z_{N}^{{}^{\prime}K}\}$ are generated by substituting $s_{i}^{j}$ into the the Geodesic curve of the category $i$. These features and the original features from the sample image, constitute a set of expanded features $\{z_{i}^{j},z_{i}^{{}^{\prime}k}|i\in(1,N),j\in(1,M_{i}),k\in(1,K)\}$. Most machine learning models can be enhanced with the expanded features.
@@ 3.4 Design of the loss function
A good loss function can improve the performance of the machine learning model. Generally, the cross entropy loss function is chosen since it can more sensitively reflect the differences between the predicted and true categories. In general, the number of generated features with the FAGC module is far more than the number of the original features. To balance the influence of the quantitative differences, a new loss function is proposed based on the cross entropy loss. we employ the standard cross entropy loss function, denoted as $L_{r}$, to evaluate the model’s performance on original features. For the generated features, two key control factors are introduced, i.e., the random probability factor $P_{g}$ and the influence factor $\lambda$. In each training batch, the final classifier loss function $L_{cls}$ is defined as follows:
@@
If $P_{g}>0.5$, the model is trained with the batch of the generated features. Otherwise, $P_{g}\leq 0$ and the generated features are ignored. $P_{g}$ can be used to randomly choose partially generated features for the training process. $\lambda$ can adjust the contribution of the generated features in the overall loss function. Thus, an influence balance between the original features and the generated features is controlled with $P_{g}$ and $\lambda$.
Figure 3: The process of the FAGC method. The hyper-sphere is a pre-shape space; the blue solid dots indicate the features in the pre-shape space; the solid orange dots represent features for parameter $v_{i}^{*}$; the solid green dots represent the two farthest known features $w_{i}^{0}$ and $w_{i}^{1}$ from features $v_{i}^{*}$; (a) Projection of image features to the pre-shape space; (b) Initialization of parameter $v_{i}^{*}$. From (c) to (g), the process of finding $w_{i}^{*}$; From (h) to (k), the process of finding new $v_{i}^{*}$; (l) Obtaining an optimal Geodesic curve $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$.
Algorithm 1 Feature Augmentation on Geodesic Curve(FAGC)
Input: Features $Q_{i}=\{z_{i}^{1},...,z_{i}^{M_{i}}\}$, number of samples in $i$-th category $M_{i}$, number of generating features $K$, and number of temporary features on the Geodesic curve $S$.
Output: Geodesic curves $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$ and the generated features $Z^{\prime}=\{z_{i}^{{}^{\prime}1},...,z_{i}^{{}^{\prime}K}\}$.
1:  $v_{i}^{*}=\mathop{\arg\max}\limits_{z_{i}^{k}}\sum_{j=1}^{M_{i}}d(z_{i}^{k},z_{i}^{j}),k\in\{1,...,M_{i}\}$;
@@
4:     $\hat{w}_{i}^{\{1,...,S\}}=\{\Gamma_{(w^{0},w^{1})}(s\cdot 1),$ $\qquad\Gamma_{(w^{0},w^{1})}(s\cdot 2),...,\Gamma_{(w^{0},w^{1})}(s\cdot S)\}$;
5:     $w_{i}^{*}=\mathop{\arg\min}\limits_{\hat{w}_{i}^{k}}\sum_{j=1}^{M_{i}}(d(z_{i}^{j},\Gamma_{(v_{i}^{*},\hat{w}_{i}^{k})}(\cdot)))^{2},$ $\qquad k\in\{1,...,S\}$;
6:     Assign $w_{i}^{*}$ to $v_{i}^{*}$;
7:  end while
8:  Calculate $d(v_{i}^{*},w_{i}^{*})$ with Formula(2);
9:  for $k=1$ to $K$ do
10:     Sample $\hat{s}_{i}^{k}\sim U(0,d(v_{i}^{*},w_{i}^{*}))$;
11:     $\hat{z}_{i}^{k}=\cos(\hat{s}_{i}^{k})\cdot v_{i}^{*}+\sin(\hat{s}_{i}^{k})\frac{w_{i}^{*}-v_{i}^{*}\cdot\cos\theta_{(v_{i}^{*},w_{i}^{*})}}{\sin\theta_{(v_{i}^{*},w_{i}^{*})}}$;
12:  end for
@@ 4 Experiments
Two datasets are utilized in the subsequent experiments. First, the proposed FAGC module is combined with various advanced learning methods and the process results demonstrate the FAGC module’s effectiveness. Then, the proposed method is compared with some data augmentation methods. Finally, some ablation experiments about classifiers are executed to check the influence of the number of features generated by FAGC. The two datasets are CIFAR-10 [45] and CIFAR-100 [45]. The benchmark methods include CoMatch [46], $\Pi$-Model [47], Mean Teacher [48], Pseudo Labeling [49], MixMatch [50], VAT [51], Dash [52], CRMatch [53], UDA [54], FixMatch [55], FlexMatch [56], FreeMatch [57], AdaMatch [58], SimMatch [59], SoftMatch [60].
4.1 Efficiently performing FAGC

### v4

@@ 3.2 Projection of the image features into the pre-shape space
According to the theory of shape space, the vertices of a shape structure are typically required to be two-dimensional or three-dimensional. Therefore, it is necessary to expand the dimensions of the features extracted from the ViT network before projecting them into the pre-shape space or shape space. The up-dimensioned step can be achieved by adding additional dimensions or using techniques such as interpolation or transformation to match the required dimensional space of the shape structure.
[…]
3.3 Construction of the Geodesic curves to fit the feature vectors
To construct a Geodesic curve in the pre-shape space, more than two image samples are required. In this paper, the number of Geodesic curves matches the number of categories, meaning that the construction of the Geodesic curves is performed within the feature vectors of the same category. The feature vectors within the same category exhibit smaller variations, making it easier to define the labels for newly augmented feature vectors. As mentioned in the above subsections, the features are extracted from the multiple images with the ViT network model. Then, the features are projected into the pre-shape space, denoted as vectors $Z=\{z_{1}^{1},...,z_{1}^{M_{1}},...,z_{N}^{1},...,z_{N}^{M_{N}}\}$, where $N$ represents the total number of image categories, $M_{i}$ represents the number of images in the $i$-th category, and $z_{i}^{j}$ denotes $j$-th sample of the $i$-th category.
The set of the Geodesic curves corresponding to different image categories in the pre-shape space is denoted as $\{\Gamma_{(v_{i}^{*},w_{i}^{*})}|i\in(1,N)\}$, where $v_{i}^{*}$ and $w_{i}^{*}$ represents the endpoints of the optimal Geodesic curve, as $v_{1}$ and $v_{2}$ in Formula (5). To obtain the $v_{i}^{*},w_{i}^{*}$ of the optimal Geodesic curve $\Gamma_{(v_{i}^{*},w_{i}^{*})}$, the first step is to calculate the Geodesic distance among all feature vectors. Then, the feature vector with the maximum sum of distances from other feature vectors is selected as the initialization parameter $v_{i}^{*}$, defined as follows:
@@ 3.3 Construction of the Geodesic curves to fit the feature vectors
where $d(z_{i}^{k},z_{i}^{j})$ means the Geodesic distance between $z_{i}^{k}$ and $z_{i}^{j}$, calculated with Formula (2). Subsequently, $w_{i}^{*}$ is obtained through Geodesic curve $\Gamma_{(w_{i}^{0},w_{i}^{1})}(\cdot)$, which is determined by the two feature vectors $w_{i}^{0}$ and $w_{i}^{1}$ furthest from $v_{i}^{*}$. As shown in Figure 3 (c), the two feature vectors, $w_{i}^{0}$ and $w_{i}^{1}$, corresponding to the maximum two Geodesic distance from $v_{i}^{*}$, are calculated using the following formulas:
@@
where $Q_{i}=\{z_{i}^{1},z_{i}^{2},...,z_{i}^{M_{i}}\}$ to represent the set of feature vectors for the $i$-th category $\{z_{i}^{1},z_{i}^{2},...,z_{i}^{M_{i}}\}$. $Q_{i}-\{\#\}$ means if the feature vectors #\# are in $Q_{i}$, they are removed from $Q_{i}$.
Then, a Geodesic curve $\Gamma_{(w_{i}^{0},w_{i}^{1})}(\cdot)$ between $w_{i}^{0}$ and $w_{i}^{1}$ is calculated with Formula (5). A set of feature vectors $\hat{w}_{i}^{\{1,...,S\}}$ is obtained by following the curve $\Gamma_{(w_{i}^{0},w_{i}^{1})}(\cdot)$, denoted as $\{\Gamma_{(w_{i}^{0},w_{i}^{1})}(s_{i}^{1}),\Gamma_{(w_{i}^{0},w_{i}^{1})}(s_{i}^{2}),...,\Gamma_{(w_{i}^{0},w_{i}^{1})}(s_{i}^{S})\}$, where $S$ represents the number of temporary feature vectors on $\Gamma_{(w_{i}^{0},w_{i}^{1})}(\cdot)$. Finally, among $v_{i}^{*}$ and $\hat{w}_{i}^{\{1,...,S\}}$, several Geodesic curves $\Gamma_{(v_{i}^{*},\hat{w}_{i}^{\{1,...,S\}})}(\cdot)$ are described. The sum of squared Geodesic distance is calculated from the $i$-th category of feature vectors $z_{i}^{1},z_{i}^{2},...,z_{i}^{M_{i}}$ to $\Gamma_{(v_{i}^{*},\hat{w}_{i}^{\{1,...,S\}})}(\cdot)$. The feature vectors $\hat{w}_{i}^{*}$ corresponding to the shortest sum is calculated as follows:
@@
and the step is illustrated in Figure 3 (f) and Figure 3 (g). Subsequently, assign the variable $w_{i}^{*}$ to $v_{i}^{*}$ and calculate the two furthest feature vectors from $v_{i}^{*}$, as illustrated in Figure 3 (h). Repeat the steps for obtaining $w_{i}^{*}$ as previously described in Figures 3 (i) to (j) until $w_{i}^{*}$ and $v_{i}^{*}$ cease to change. Consequently, $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$ denotes the optimal Geodesic curve fitting the distribution of feature vector $z_{i}^{*}$, as depicted in Figure 3 (k). The detailed process of the FAGC processing specific category feature vectors is outlined in Algorithm  1.
All Geodesic curves for different categories, denoted as $\{\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)|i\in\{1,N\}\}$, can be constructed with the above mentioned method. And, with utilizing Formula (2), the Geodesic distance $d(v_{i}^{*},w_{i}^{*})$ between all pairs of $v_{i}^{*}$ and $w_{i}^{*}$ is calculated. The distance indicates the range of the angle value $s$ within the Geodesic curve function $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$. A uniform distribution $s\sim U(0,d(v_{i}^{*},w_{i}^{*}))$ is employed to sample a set of angle values $\{s_{i}^{j}|j\in\{1,K\}\}$, where $K$ represents the number of sampled angle values for the Geodesic curve of the $i$-th category. The new feature vectors $Z^{\prime}=\{z_{1}^{{}^{\prime}1},...,z_{1}^{{}^{\prime}K},...,z_{N}^{{}^{\prime}1},...,z_{N}^{{}^{\prime}K}\}$ are generated by substituting $s_{i}^{j}$ into the the Geodesic curve of the category $i$. These feature vectors and the original feature vectors from the sample image, constitute a set of augmented feature vectors $\{z_{i}^{j},z_{i}^{{}^{\prime}k}|i\in(1,N),j\in(1,M_{i}),k\in(1,K)\}$.
@@ 3.4 Design of the loss function
To better balance the weight of original and augmented feature vectors during the training process, we optimized the loss function for the application of the FAGC. Generally, the cross entropy loss function is chosen since it can more sensitively reflect the differences between the predicted and true categories. In general, the number of augmented feature vectors with the FAGC is far more than the number of the original feature vectors. To balance the influence of the quantitative differences, a new loss function is proposed based on the cross entropy loss. We employ the standard cross entropy loss function, denoted as $L_{r}$, to evaluate the model’s performance on original feature vectors. For the augmented feature vectors, two key control factors are introduced, i.e., the random probability factor $P_{g}$ and the influence factor $\lambda$. Let $\sigma$ be a random variable uniformly distributed in $[0,1]$, then the $P_{g}$ can be computed as
$P_{g}=\begin{cases}1&\text{if }\sigma>0.5\\ 0&\text{otherwise}\end{cases}.$ (1)
In each training batch, the final classifier loss function $L_{cls}$ is defined as follows:
@@
If $P_{g}=1$, the model is trained with the batch of the augmented feature vectors. Otherwise, $P_{g}=0$ and the augmented feature vectors are ignored. $P_{g}$ can be used to randomly choose partially augmented feature vectors for the training process. $\lambda$ adjusts the contribution of the augmented feature vectors in the overall loss function and takes values between 0 and 1. Thus, an influence balance between the original feature vectors and the augmented feature vectors is controlled with $P_{g}$ and $\lambda$.
Figure 3: The process of constructing an optimal Geodesic curve and generating new feature vectors using the FAGC. The hyper-sphere is a pre-shape space; the blue solid dots indicate the feature vectors in the pre-shape space; the solid orange dots represent feature vectors for parameter $v_{i}^{*}$; the solid green dots represent the two farthest known feature vectors $w_{i}^{0}$ and $w_{i}^{1}$ from feature vector $v_{i}^{*}$. (a) Projection of image features to the pre-shape space; (b) initialization of parameter $v_{i}^{*}$; from (c) to (g), the process of finding $w_{i}^{*}$; from (h) to (k), the process of finding new $v_{i}^{*}$; (l) obtaining an optimal Geodesic curve $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$ and augmented feature vectors.
Algorithm 1 Feature Augmentation on the Geodesic Curve in the Pre-Shape Space (FAGC)
Input: Feature vectors $Q_{i}=\{z_{i}^{1},...,z_{i}^{M_{i}}\}$ in the pre-shape space, number of samples in $i$-th category $M_{i}$, number of augmented feature vectors $K$, and number of temporary feature vectors on the Geodesic curve $S$.
Output: The Geodesic curve $\Gamma_{(v_{i}^{*},w_{i}^{*})}(\cdot)$ and the augmented feature vectors $Z^{\prime}=\{z_{i}^{{}^{\prime}1},...,z_{i}^{{}^{\prime}K}\}$.
[…]
@@
4:  $w^{1}=\mathop{\arg\max}\limits_{{Q_{i}-\{v_{i}^{*},w^{0}\}}}\left(d(v_{i}^{*},{Q_{i}-\{v_{i}^{*},w^{0}\}})\right)$;
5:  $\hat{w}_{i}^{\{1,...,S\}}=\{\Gamma_{(w^{0},w^{1})}(s\cdot 1),$ $\qquad\Gamma_{(w^{0},w^{1})}(s\cdot 2),...,\Gamma_{(w^{0},w^{1})}(s\cdot S)\}$;
6:  $w_{i}^{*}=\mathop{\arg\min}\limits_{\hat{w}_{i}^{k}}\sum_{j=1}^{M_{i}}(d(z_{i}^{j},\Gamma_{(v_{i}^{*},\hat{w}_{i}^{k})}(\cdot)))^{2},$ $\qquad k\in\{1,...,S\}$;
7:  Assign $w_{i}^{*}$ to $v_{i}^{*}$;
8:end while
9: Calculate $d(v_{i}^{*},w_{i}^{*})$ with Formula(2);
10:for $k=1$ to $K$ do
11:  Sample $\hat{s}_{i}^{k}\sim U(0,d(v_{i}^{*},w_{i}^{*}))$;
12:  $\hat{z}_{i}^{k}=\cos(\hat{s}_{i}^{k})\cdot v_{i}^{*}+\sin(\hat{s}_{i}^{k})\frac{w_{i}^{*}-v_{i}^{*}\cdot\cos\theta_{(v_{i}^{*},w_{i}^{*})}}{\sin\theta_{(v_{i}^{*},w_{i}^{*})}}$;
13:end for
@@ 4 Experiments
In this section, the FAGC is validated across several commonly small-sample scenarios, encompassing different learning frameworks and datasets of varying scales and difficulties.
4.1 Experimental settings
The FAGC framework operates in an offline manner and is structured into three distinct phases: feature extraction, feature augmentation, and downstream task integration. Since the performance of the FAGC depends on the quality of the features, the ViT [44] series networks, primarily ViT-S [44] and ViT-T [44], are used as feature extractors. Except for Section 4.4, the experimental data undergo a preprocessing process that includes horizontal flips and random crops before feature extraction. The standard cross-entropy loss function is employed, with AdamW [45] as the optimizer. Additional configurations include a weight decay of 0.0005, a learning rate of 0.0005, a layer decay rate of 0.5, and a moving average momentum of 0.9. The $\lambda$ is set to 0.5. All experiments are run on a server equipped with an NVIDIA 4090 GPU, and the reported results represent the average of three experimental runs. Other settings related to specific experimental content will be explained in subsequent sections.
In this paper, the FAGC is first integrated into the semi-supervised learning paradigm. Since real-world applications often face a scarcity of data labels, semi-supervised learning represents one of the most common approaches in deep learning. Unlike traditional machine learning pipelines, which is inherently compatible with the FAGC’s offline training, applying the FAGC to deep learning requires certain adaptations. Specially, when integrating the FAGC into a target deep model, the architecture of model is modularized into two core components: a feature extractor and a classifier. In practice, the model is first trained on the real data using conventional deep learning methods, and the trained model is retained. Subsequently, features are extracted from the input data and passed through the FAGC for augmentation. The augmented feature vectors, combined with the original feature vectors, are then utilized to fine-tune the classifier, thereby enhancing the model’s overall performance.
4.2 Applying the FAGC in existing semi-supervised or supervised methods